OpenStudy (anonymous):

Factor the polynomial: 54k^2-24

OpenStudy (anonymous):

get common factor of 54 and 24.

OpenStudy (anonymous):

so they are both even numbers, so lets start factoring 2. 54k^2-24 = 2(27)k^2-2(12)= 2(27k^2-12)

OpenStudy (anonymous):

what can you do next? do you see that 27 and 12 are both divisible by 3? so factor out the three. 2(27k^2-12)= 2((3)9k^2-3(4))=6(9k^2-4)

OpenStudy (anonymous):

I'm sorry you lost me

OpenStudy (anonymous):

she factored 6 out of 54k^2-24 making it 6*(9k^2-4) factoring 9k^2-4 is easier than 54k^2-24

OpenStudy (anonymous):

@robz8 and after that, what do I do?

OpenStudy (anonymous):

factor 9k^2-4